OpenStudy (anonymous):

Kelly and Daniel wrote the following proofs to prove that vertical angles are congruent. Who is correct? Both Kelly and Daniel are correct. Neither Kelly or Daniel is correct. Kelly is correct, but Daniel is not. Daniel is correct, but Kelly is not.

OpenStudy (anonymous):

Daniel’s Proof Statement Justification 1 + 2 = 180° Definition of Supplementary Angles 1 + 4 = 180° Definition of Supplementary Angles 1 + 2 = 1 + 4 Transitive Property of Equality 2 = 4 Subtraction Property of Equality

OpenStudy (anonymous):

User: Brooke Dalton In Course: Geometry V14 ( 3801) Instructor: Ricardo Roque 09.03 Line Proofs Warning: There is a checkbox at the bottom of the exam form that you MUST check prior to submitting this exam. Failure to do so may cause your work to be lost. Question 1 (Multiple Choice Worth 4 points) Kelly and Daniel wrote the following proofs to prove that vertical angles are congruent. Who is correct? Kelly’s Proof Statement Justification 2 = 4 Vertical angles are congruent 1 = 3 Vertical angles are congruent Vertical Angles are congruent. Vertical Angle Theorem

OpenStudy (anonymous):

i am on that same problem...lol i need help too